A trapezoid has three sides of length 4 inches and one side of length 8 inches. What is its area?
Answer by rfer(2688)

find the heigth
16-4=12
sq rt of 12=3.46
------------------
A=h((b1+b2)/2)
A=3.46(12/2)
A=3.46*6
A=20.76 sq in